iYuko (iyuko):

Introduction to Factoring Lesson 1 Come learn with me!

OpenStudy (anonymous):

hey rational

OpenStudy (rational):

heyy

iYuko (iyuko):

Recall that when 2 or more numbers, variables, or algebraic expressions are multiplied, each is called a factor When you are asked to factor a number or a algebraic expression, you are being asked, "What, when multiplied, will give you that number or expression?" For example, you can factor 6 as 3*2 since 3*2=6. You can factor 15x^5 as 3x^2 * 5x^3 since 3x^2 * 5x^3 =15x^5. Factoring is simply the reverse of multiplying. 6 and 15x^5 are simple expressions to factor and can be factored in many ways. The factors of the polynomial \[2x ^{3} +X-12\] are not east to recognize In this tutorial we will be learning the techniques to find the factors of a polynomial. We will begin with simple factoring.

iYuko (iyuko):

When we factor, we are looking for the largest common factor. For example, in the polynomial 48x-16y, a common factor is 2. We could factor 48-16y as 2(24-8y). However, this is not complete. To factor 48-16y completely, we look for the greatest common factor of 48 and 16. 48-16y= 16(3x-y)

iYuko (iyuko):

Common factors of a polynomial 1. You can determine the numerical common factor by asking, "What is the largest integer that will divide into the coefficient of each term?" 2. You can determine the variable common factor by asking, "What variables are common to each term, and what is the largest exponent of those variables that is common?"

iYuko (iyuko):

Lets try an example: Factor. \[8x ^{3}y+16x ^{2}y ^{2}+24x ^{3}y ^{3}\] We see that 8 is the largest integer that will divide equally into the 3 numerical coefficients. We can factor \[x ^{2}\] out of each term (not just x). For the other variable we factor y out of each term. \[8x ^{3}y+16x ^{2}y ^{2}+24x ^{3}y ^{3}=8x^{2}y(x+2y+3yx ^{2})\] To check this you multiply the factors you obtain. The result should be the same as the original polynomial.

iYuko (iyuko):

Lesson 2 (Factor by grouping) should be up sometime tomorrow.

iYuko (iyuko):

Let me know of some topics you would like me to do. :O \(\hspace{0.7pt}{\rlap{\Huge\cal\color{red}{\bigstar}\color{orange}{i}\color{yellow}{Y}\color{green}{u}\color{blue}{k}\color{indigo}{o}\color{violet}{\bigstar}}}\)

OpenStudy (girllover123):

@logan13

iYuko (iyuko):

Change In plans, lesson 2 will not be out today sorry for any inconvenience. Im shooting for Friday. -iYuko

OpenStudy (anonymous):

GOOD JOB, @iYuko!

OpenStudy (anonymous):

i want a medallllll T_T

OpenStudy (anonymous):

I gave you a medal because I love the hunger games

OpenStudy (anonymous):

XD thx :D and ikr? they're the best